# Adler function, Bjorken polarized Sum Rule – confirmation of elements of the $\{\beta\}$-expansion

### v1

@@ Appendix A QCD $\beta$-function
$\beta_{1}$ $=$ $-4\,\underline{C_{\text{F}}T_{r}n_{f}}+\frac{34}{3}\,C_{\text{A}}^{2}+\frac{20}{3}\,C_{\text{A}}T_{r}n_{f};$ (21b)
@@
$\beta_{3}$ $=$ $\frac{1232}{243}\,\underline{C_{F}(T_{R}n_{f})^{3}}+\underline{(C_{F}T_{R}n_{f})^{2}}\left(\frac{1352}{27}-\frac{704}{9}\zeta_{3}\right)+46\underline{(C_{F})^{3}T_{R}n_{f}}+C_{A}^{3}T_{R}n_{f}\left(-\frac{39143}{81}+\frac{136}{3}\zeta_{3}\right)$ (21e)
@@
$T_{R}=\frac{1}{2}\,;\;C_{\text{F}}=\frac{N^{2}-1}{2N}\,;\;C_{A}=N\,;\;N_{A}=2C_{\text{F}}C_{\text{A}}\equiv N^{2}-1\,;$
@@
Appendix B Numerical estimates of the elements $d_{3,4}[.],c_{3,4}[.]$ of $\{\beta\}$-expansion
We present the numerical estimates for the each of elements, discussed in Sec.4 for reader’s convenience. For the expressions $d_{3}[.]$ in Eq.(4) and similar ones $c_{3}[.]$ Baikov:2022zvq we have Kataev:2014zha ; Kataev:2014zwa
$\!\!\!\!{\color[rgb]{1,0,0}d_{3}[2]}$$\approx$ ${\color[rgb]{1,0,0}\bm{12.41}},~{}~{}{\color[rgb]{1,0,0}d_{3}[0,1]\approx\bm{-4.799}},d_{3}[1]\approx 293.55,~{}~{}d_{3}[0]\approx-2295.84\,,$ (24a)
$\!\!\!\!{\color[rgb]{1,0,0}c_{3}[2]}$$\approx$ ${\color[rgb]{1,0,0}\bm{-25.56}},~{}{\color[rgb]{1,0,0}c_{3}[0,1]\approx\bm{-0.43}},c_{3}[1]\approx-159.84,~{}c_{3}[0]\approx 2242.51\,.$ (24b)
@@
${\color[rgb]{1,0,0}d_{4}[2]}$$\approx$ $154.47,~{}d_{4}[0,1]\approx 195.46,~{}d_{4}[1]\approx-550.54,$ (25b)
@@
The elements marked in red in (24,25,26) have been confirmed, at least in their abelian part, by the diagrammatical calculations in the standard QCD in Secs.2, 3. The widely debated “zero” element $d_{4}[0]$ reads $-98.069+0.0802\,n_{f}$ in the case of the using the coupling constant $\alpha_{s}/\pi$ instead of $a_{s}$.

### v3

@@ Appendix A QCD $\beta$-function
$\beta_{1}$ $=$ $-4\,\underline{C_{\text{F}}T_{R}n_{f}}+\frac{34}{3}\,C_{\text{A}}^{2}-\frac{20}{3}\,C_{\text{A}}T_{R}n_{f};$ (24b)
@@
$\beta_{3}$ $=$ $\frac{1232}{243}\,\underline{C_{F}(T_{R}n_{f})^{3}}+\underline{(C_{F}T_{R}n_{f})^{2}}\left(\frac{1352}{27}-\frac{704}{9}\zeta_{3}\right)+46\underline{(C_{F})^{3}T_{R}n_{f}}+C_{A}^{3}T_{R}n_{f}\left(-\frac{39143}{81}+\frac{136}{3}\zeta_{3}\right)$ (24d)
@@
$T_{R}=\frac{1}{2}\,;\;C_{\text{F}}=T_{R}\frac{N^{2}-1}{N}\,;\;C_{A}=N\,;\;N_{A}=2C_{\text{F}}C_{\text{A}}\equiv N^{2}-1\,;$ (25)
@@
Appendix B Numerical estimates of the elements $d_{2,3,4}[.],c_{2,3,4}[.]$ of $\{\beta\}$-expansion
We present the numerical estimates for the each of elements, discussed in Sec.4 for reader’s convenience. For the expressions $d_{2,3}[.]$ in Eq.(3,4) and similar ones $c_{2,3}[.]$ Baikov:2022zvq we have Kataev:2014zha ; Kataev:2014zwa ,
$\!\!\!\!{\color[rgb]{1,0,0}d_{2}[1]}$$\approx$ ${\color[rgb]{1,0,0}\bm{2.77}},~{}~{}~{}~{}~{}{\color[rgb]{1,0,0}d_{2}[0]}={\color[rgb]{1,0,0}\bm{4/3}},~{}~{}~{}~{}~{}~{}~{}{\color[rgb]{1,0,0}c_{2}[1]}={\color[rgb]{1,0,0}\bm{-8}},~{}~{}~{}~{}~{}~{}{\color[rgb]{1,0,0}c_{2}[0]}={\color[rgb]{1,0,0}\bm{44/3}};$ (28a)
$\!\!\!\!{\color[rgb]{1,0,0}d_{3}[2]}$$\approx$ ${\color[rgb]{1,0,0}\bm{12.41}},~{}~{}{\color[rgb]{1,0,0}d_{3}[0,1]}\approx{\color[rgb]{1,0,0}\bm{-4.799}},d_{3}[1]\approx 293.55,~{}~{}d_{3}[0]\approx-2295.84\,,$ (28b)
$\!\!\!\!{\color[rgb]{1,0,0}c_{3}[2]}$$\approx$ ${\color[rgb]{1,0,0}\bm{-25.56}},~{}{\color[rgb]{1,0,0}c_{3}[0,1]\approx\bm{-0.43}},c_{3}[1]\approx-159.84,~{}c_{3}[0]\approx 2242.51\,.$ (28c)
@@
${\color[rgb]{1,0,0}d_{4}[2]}$$\approx$ ${\color[rgb]{1,0,0}\bm{154.47}},~{}d_{4}[0,1]\approx 195.46,~{}d_{4}[1]\approx-550.54,$ (29b)
@@
The elements marked in red in (28,29,30) have been confirmed, at least in their abelian part, by the diagrammatical calculations in the standard QCD in Secs.2, 3. The widely debated “zero” element $d_{4}[0]$ reads $-98.069+0.0802\,n_{f}$ in the case of the using the coupling constant $\alpha_{s}/\pi$ instead of $a_{s}$.